I've been fighting a switch matrix issue with a Who Dunnit and am ready to throw in the towel; this one has me stumped.
At certain points during the game the software will dim general illumination and start all three reels spinning, and will then stop the reels one at a time from left to right; the DMD animation shows the high score initials while the reels are spinning, and as the reels stop the symbols on the reels are displayed on the DMD. OCCASIONALLY the reels don't stop, and the game software seems to be stuck in an odd state: if I manually activate any playfield switch the software "sort of" responds (e.g., it may play a different audio clip, GI comes back on briefly), but the slot reels keep spinning. The software doesn't enter ball search mode or give any indication it thinks anything is out of order.
I've found that pulling the connector from the left reel index opto "breaks" the game software out of this odd state, the DMD animation changes to display the (now stopped) reel symbols, and things go back to normal. I've replaced the index opto (no change), and have carefully examined all switches on the row/column of the left reel index opto for issues, but nothing seems amiss. The reel test diagnostic passes with no issues, and 99% of the time the interaction with the slot machine works just fine. If I disable the slot machine via the game settings the game software never gets into the odd state.
My assumption is the game software is waiting for a switch closure indication from the left reel index opto, and for some reason it is not getting one. The reel index optos are directly driven from the switch matrix just as with the Twilight Zone clock optos; the same part number are used in W?D and TZ.
I've tried both software versions for the game (problem still occurs), and have swapped out both the CPU and power driver boards with known good spares (problem still occurs). I also tried bizarre things like disconnecting all GI and controlled lamps and running the game with the room lighting off on the odd chance stray light leakage is confusing the left index opto receiver. This appears to be a switch matrix issue that is interfering with the left index opto, but I've run out of ideas.
Has anyone seen this issue with W?D, or have any thoughts on a possible solution (other than a stick of dynamite)? Oddball theories are most welcome at this point.